OpenStudy (anonymous):

-x^3+3x^2+24x-80 C = - 5 what are the remaining zeros? what is this in factored form ?

OpenStudy (anonymous):

x=-5 and x=4 (x+5)(-x+4)(x-4)

OpenStudy (anonymous):

(x+5) ( -x^2 +x -16) Just x = -5 is real root The other is complex roots

OpenStudy (anonymous):

^ that answer is terribly wrong

OpenStudy (anonymous):

\[-(x-4)^2 (x+5)\] zeros are 4 and -5

OpenStudy (anonymous):

OOps (x+5) ( -x^2 +8x -16)

OpenStudy (anonymous):

-(x + 5) ( x -4)^2 x = -5 and x = 4